OpenStudy (anonymous):

if g(x) = 2x - 4 and g(m) = -4 and g(n) = 6, find g(m+n) i got -4 can someone check that? i want to make sure its right!

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Ahsome

OpenStudy (ahsome):

\(g(m)=-4\) \(g(n)=6\) \(g(m+n)\) Sub the values \(-4+6\) \(2\)

OpenStudy (anonymous):

but g(x) = 2x-4 so dont you sub -4 and 6 in for x?

OpenStudy (anonymous):

@Ahsome

OpenStudy (ahsome):

Yes, you are right. I didn't notice the equation.

jimthompson5910 (jim_thompson5910):

yeah you need to find m & n first, then compute g(m+n)

OpenStudy (anonymous):

is -4 right?

OpenStudy (anonymous):

@jim_thompson5910 @Ahsome

jimthompson5910 (jim_thompson5910):

what did you get for m?

OpenStudy (anonymous):

@jim_thompson5910 for m i got -12 and for n i got -4

jimthompson5910 (jim_thompson5910):

g(x) = 2x - 4 g(m) = 2m - 4 -4 = 2m - 4 m = ??

OpenStudy (anonymous):

if you plug -4 in for m you-4 i thought because i did 2(-4)-4 which is -12

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Ahsome

OpenStudy (anonymous):

@nikato

jimthompson5910 (jim_thompson5910):

your task is to now solve -4 = 2m - 4 for m

jimthompson5910 (jim_thompson5910):

you are NOT plugging in m = -4

OpenStudy (anonymous):

0=2m?

OpenStudy (anonymous):

0/2=m?

jimthompson5910 (jim_thompson5910):

so m = 0

jimthompson5910 (jim_thompson5910):

what is the value of n?

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

m+n = ??

OpenStudy (anonymous):

5 thanks!

jimthompson5910 (jim_thompson5910):

now compute g(m+n) = g(5) = ??

OpenStudy (anonymous):

oh 6

OpenStudy (anonymous):

so the final answer is 6

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

ok thanks!

jimthompson5910 (jim_thompson5910):

np